# Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain

### v2

@@ IV-B Patch Embedding
Specifically, we first utilize a $7\times 7$ convolutional layer with a stride of 4 to model the local correlation of the input image. Then, we downsample the feature map twice through two downsampling blocks. Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer. The details of the adaptation with the two base models [1] (Vim-Ti and Vim-S) are shown in Table I.
TABLE I: Convolutional stem for patch embedding.
@@
Block1 56$\times$56 7$\times$7, 48, stride 4 7$\times$7, 48, stride 4
@@
Block2 28$\times$28 $\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
Block3 14$\times$14 $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
@@
IV-C Vim-F Block
As mentioned earlier, the Vim-F block is an improvement over the Vim block, as shown in Fig. 3. Specifically, a 2D FFT is performed on the input to obtain the spectrogram. Then, the spectrogram is added to the original input, with the intensity of the frequency and spatial domain information being adjusted by two trainable parameters, $\alpha$ and $\beta$, respectively. After the linear layer mixes the hidden dimensions of the feature map, the cross-domain information is sent to the Vim block for processing.
Figure 3: Vim-F block
@@ IV-C Vim-F Block
Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation. The Cooley-Tukey FFT algorithm decomposes a DFT of a sequence with length N into two DFTs of sub-sequences with length N/2, significantly reducing the computational complexity. During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results. For a 2D-DFT with dimensions $(L,D)$, its complexity is $\Omega(L\cdot D\cdot logL\cdot logD)$. The feedforward network is implemented by a linear layer, with a complexity of $\Omega(L\cdot D^{2})$. Therefore, compared to the Vim block, the additional overhead of the Vim-F block is approximately linearly related to the sequence length $L$.
IV-D Vim-F(H) Block
@@
Observing Eq. (1), it is evident that since all elements of matrix $A$ are strictly confined within the range of 0 to 1 [5], the early information stored in the hidden state $h$ will continuously decay. This leads to Mamba being sensitive to the order of the input sequence. We propose that by weighting the input sequence with attention mechanisms before processing with Mamba, this strong local bias can be mitigated, thereby enhancing Mamba’s capability to handle long sequences. We integrate the Vim-F block with the linear attention [11] to maintain approximately linear complexity, which we term Vim-F(H), and its structure is depicted in Fig. 4.
Figure 4: Vim-F(H) block
TABLE II: Comparison of different backbones on ImageNet-1K classification.
@@
VMamba-T [2] 22 5.6 82.2
V Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation. The specific architectures of Vim-F and Vim-F(H) are as follows: We replaced only the first 6 Vim blocks (25% of the total number of blocks) with either Vim-F or Vim-F(H) blocks to balance performance and efficiency while maintaining approximately the same number of parameters. The number of linear attention heads in Vim-F(H) is set to 24.
V-A Image Classification
@@
Settings. We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Due to hardware limitations, our experiments are performed on 3 A6000 GPUs. Therefore, we adjust the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling rule[18].
@@
Results. We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table II. Vim-F outperforms the original Vim in terms of performance, and Vim-F(H) further enhances Vim-F’s performance by introducing linear attention. It is worth noting that SOTA ViMs such as VMamba [2], PlainMamba [9], and EfficientVMamba [4] all adopt more complex CNN-Mamba hybrid encoding architectures, and their network structures are similar to advanced convolutional networks like ConvNeXt [19]. Since these works have not conducted the necessary comparative experiments to address this issue, it is reasonable to suspect that the excellent performance of these ViMs may primarily stem from the use of depthwise separable convolutions. However, Vim-F still uses a pure SSM encoder to model visual information. LocalVim[3] is one of the representative works that improve the scanning strategy for Vim. LocalVim has designed 8 scanning strategies and employs neural architecture search to determine the scanning strategy used by each of the 4 scanning branches in every LocalVim block. Additionally, it utilizes a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in each branch feature. Moreover, the total number of blocks in LocalVim has not followed Vim’s 24 but has been reduced to 20. In comparison, our works achieve a better balance between performance, efficiency, and model complexity.
TABLE III: Object detection and instance segmentation results on COCO.
@@
ResNet-18 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
Vim-F-Ti (ours) 28M 189G 37.7 60.8 40.9 35.7 57.8 38.0
ResNet-50 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
Vim-F(H)-Ti (ours) 28M 189G 38.2 61.8 41.7 35.9 58.6 38.0
EfficientVMamba-S 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
ResNet-101 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
Vim-S 44M 272G 40.9 63.9 45.1 37.9 60.8 40.7
Vim-F-S (ours) 47M 273G 41.6 64.1 45.5 38.4 60.9 41.3
Swin-T 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
Vim-F(H)-S (ours) 57M 275G 42.9 65.4 47.1 39.2 62.1 42.1
Swin-S 69M 354G 44.8 66.6 48.9 40.9 63.2 44.2
V-B Object Detection and Instance Segmentation
Settings. We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F for object detection and instance segmentation on the MSCOCO 2017 dataset. Since the backbone of Vim-F is non-hierarchical, following ViTDet[20], we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector. The remaining settings were consistent with Swin[17]. Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule). The learning rate is initialized at $1\times 10^{-4}$ and is reduced by a factor of 10×10$\times$10 × at the 9th and 11th epochs.
@@ V-B Object Detection and Instance Segmentation
Results. We summarize the comparison results of Vim-F and Vim-F(H) with other backbones in Table III. It can be seen that our Vim-F consistently outperforms Vim. Similar to the results in classification tasks, Vim-F(H) exhibits better performance than Vim-F and is competitive with advanced ViT architectures.
VI Conclusion
@@
One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data. Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships. We observe that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field. We adapt a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains. Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps. We also introduce additional local correlations during patch embedding by using overlapping convolutions. Vim-F(H) enhances the performance of Vim-F by introducing linear attention, which mitigates the sequence order sensitivity of the Mamba encoder. Extensive experiments demonstrate that our work effectively improves the performance of Vim, with Vim-F and Vim-F(H) achieving competitive results.

### v3

@@ 4.2 Patch Embedding
Specifically, we first utilize a $7\times 7$ convolutional layer with a stride of 4 to model the local correlation of the input image. Then, we downsample the feature map twice through two downsampling blocks. Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer. The details of the adaptation with the two base models [22] (Vim-Ti and Vim-S) are shown in Table 1.
Table 1: Convolutional stem for patch embedding.
@@
Block1
$\times56$
$\times7$, 48, stride 4 $\times7$, 48, stride 4
@@
Block2
$\times28$
$\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
Block3
$\times14$
$\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
@@
4.3 Vim-F Block
@@
As mentioned earlier, the Vim-F block is an improvement over the Vim [22] block, as shown in Fig. 3(a). Specifically, a 2D FFT is performed on the input to obtain the spectrogram. Then, the spectrogram is added to the original input, with the intensity of the frequency and spatial domain information being adjusted by two trainable parameters, $\alpha$ and $\beta$, respectively. After the linear layer mixes the hidden dimensions of the feature map, the cross-domain information is sent to the Vim block for processing. Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation. The Cooley-Tukey FFT algorithm decomposes a DFT of a sequence with length $N$ into two DFTs of sub-sequences with length $N/2$, significantly reducing the computational complexity. During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results. For a 2D-DFT with dimensions $(L,D)$, its complexity is $\Omega(L\cdot D\cdot logL\cdot logD)$. The complexity of the linear layer is $\Omega(L\cdot D^{2})$. Therefore, compared to the Vim block, the additional overhead of the Vim-F block is approximately linearly related to the sequence length $L$. Considering that Vim [22] flattens the input into a sequence of length 196, the hidden state dimensions $D$ of the two variants Vim-Ti and Vim-S are 192 and 384 respectively, with a state expansion ratio of 2. These additional computational overheads are even significantly smaller than any single linear projection in Eq. (4).
Figure 3: The architecture of the Vim-F block and the Vim-F(H) block.
4.4 Vim-F(H) Block
@@
Observing Eq. (2), it is evident that since all elements of matrix $A$ are strictly confined within the range of 0 to 1 [2], the early information stored in the hidden state $h$ will continuously decay. This leads to Mamba being sensitive to the order of the input sequence. We propose that by weighting the input sequence with attention mechanisms before processing with Mamba, this strong local bias can be mitigated, thereby enhancing Mamba’s capability to handle long sequences. We integrate the Vim-F block with the linear attention [6] to maintain approximately linear complexity, which we term Vim-F(H), and its structure is depicted in Fig. 3(b).
Figure 4: Illustration of (a) Overall architecture of Vim-F(CF), (b) Vim block structure for reference, and (c) Proposed Vim-F(CF) block structure.
4.5 Vim-F(CF): Is the convolution necessary in the ViM encoder?
We know that Mamba [2] was originally designed for NLP, with its core module S6 being good at capturing long-range dependencies. Convolutions were introduced to compensate for the model’s limitations in short-range context modeling. However, the sequence lengths processed by ViMs are usually short. For example, in Vim [22] and PlainMamba [20], images are flattened into sequences with a length of 196. Thus, from this perspective, retaining convolutions in the ViM encoder is unnecessary. Due to the influence of traditional visual model design concepts and experimental results, the majority of existing ViM encoders retain convolutional layers. VMamba [9] points out that depthwise convolutions play an important role in forming a global effective receptive field (ERF). We believe that removing the convolutional layer and adopting frequency domain scanning can make the modeling approach of ViM closer to that of Transformer, thereby enhancing the scalability and transferability of ViM.
We use convolution-free Vim-F blocks to build the proof-of-concept network Vim-F(CF), as shown in Fig. 4. Vim-F(CF) follows a classic 4-stage structure to enable SSM to model visual features across different scales. Unlike common CNNs, Transformers, ViMs, and their hybrid models, Vim-F(CF) does not incorporate attention mechanisms or feed forward network (FFN) modules.
5 Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection, instance segmentation, and semantic segmentation. The specific architectures of Vim-F and Vim-F(H) are as follows: We first replace the patch embedding module, and then only the first 6 Vim blocks (25% of the total number of blocks) are substituted with either Vim-F or Vim-F(H) blocks to balance performance and efficiency. The number of linear attention heads in Vim-F(H) is set to 24.
Table 2: Comparison of different backbones on ImageNet-1K classification.
@@ 5 Experiments
VMamba-T [9] 30 4.9 82.6
5.1 Image Classification
5.1.1 Settings.
@@ 5.1 Image Classification
We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Due to hardware limitations, our experiments are performed on 3 A6000 GPUs. Therefore, we adjust the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling rule[1].
5.1.2 Results.
@@
We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table 2. Vim-F outperforms the original Vim in terms of performance, and Vim-F(H) further enhances Vim-F’s performance by introducing linear attention. Although not meticulously designed, the proof-of-concept network Vim-F(CF) matches Vim-S in performance and computational complexity and has a parameter count advantage, indicating that convolution-free ViM encoders and associated model architectures are also a promising research direction. It is worth noting that SOTA ViMs such as VMamba [9], PlainMamba [20], and EfficientVMamba [14] all adopt more complex CNN-Mamba hybrid encoders, using depthwise separable convolutions to extract features in a 2D space. In addition, their network architectures also draw on advanced networks like ConvNeXt [11], MetaFormer [21], MobileViT [13]. Since these works have not conducted ablation experiments on the necessity of the SSM block, it is reasonable to suspect that the exceptional performance of these ViMs may mainly be attributed to the introduction of depthwise separable convolutions and advanced network architectures. Unlike these works, Vim-F and Vim-F(H) do not use depthwise separable convolutions and largely follow the design philosophy of Vim’s network architecture. LocalVim[5] is one of the representative works that improve the scanning strategy for Vim. LocalVim has designed 8 scanning strategies and employs neural architecture search to determine the scanning strategy used by each of the 4 scanning branches in every LocalVim block. Additionally, it utilizes a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in each branch feature. Moreover, the total number of blocks in LocalVim has not followed Vim’s 24 but has been reduced to 20. In comparison, our works achieve a better balance between performance, efficiency, and model complexity.
Table 3: Ablation study on the proposed patch embedding module and Vim-F block proportion.
Method Params (M) FLOPs (G) Top-1 (%)
Vim-Ti [22] 7 1.5 76.1
- Position Embedding 7 1.5 74.8
Vim-Ti-F-V0 7 1.5 76.3
Vim-Ti-F-V1 7 1.5 76.7
Vim-Ti-F-V2 8 1.7 76.8
Vim-Ti-F-V3 9 1.9 77.0
5.1.3 Ablation Study on ImageNet-1K.
Here, taking Vim-Ti-F as an example, we conduct an ablation study on the effectiveness of our proposed patch embedding module and the impact of the Vim-F block proportion. We configured four variants of Vim-Ti-F, namely V0 to V3, with Vim-F block proportions set at 0%, 25%, 50%, and 100%, respectively. As shown in Table 3, removing position embedding caused a 1.3% performance degradation of Vim-Ti. Vim-Ti-F-V0 showed slightly better performance than Vim-Ti, indicating that using a convolutional stem with overlapping convolutions as patch embedding is a more effective approach. Additionally, substituting more Vim blocks with Vim-F blocks can lead to steady performance improvement. However, due to the overall representation limitations of the model, it is unnecessary to use Vim-F blocks exclusively.
Table 4: Object detection and instance segmentation results on COCO.
Backbone Params FLOPs APb $^{\mathrm{b}}_{50} AP$$^{\mathrm{b}}_{75}$ APm $^{\mathrm{m}}_{50} AP$$^{\mathrm{m}}_{75}$
ResNet-18 [4] 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
Vim-Ti [22] 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
Vim-Ti-F (ours) 28M 189G 37.7 60.8 40.9 35.7 57.8 38.0
ResNet-50 [4] 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
Vim-Ti-F(H) (ours) 28M 189G 39.3 61.8 42.6 36.7 58.9 39.2
EfficientVMamba-S [14] 31M 197G 39.5 60.8 43.1 35.8 57.4 38.3
ResNet-101 [4] 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
Vim-S [22] 44M 272G 40.9 63.9 45.1 37.5 59.4 40.2
Vim-S-F (ours) 47M 273G 41.6 62.7 45.6 37.9 60.8 40.7
Vim-F(CF) (ours) 42M 272G 42.0 62.5 45.9 38.4 60.9 41.3
Swin-T [10] 48M 267G 42.7 65.2 46.8 39.2 62.1 42.1
Vim-S-F(H) (ours) 57M 275G 42.9 65.4 47.1 39.3 62.2 42.2
5.2 Object Detection and Instance Segmentation
5.2.1 Settings.
@@ 5.2 Object Detection and Instance Segmentation
5.2.2 Results.
@@
We summarize the comparison results of Vim-F, Vim-F(H), and Vim-F(CF) with other backbones in Table 4. It can be seen that our Vim-F consistently outperforms Vim. Notably, Vim-S-F outperforms Vim-F(CF) in classification tasks, while Vim-F(CF) surpasses Vim-S-F in downstream dense prediction tasks. Similar to the results in classification tasks, Vim-F(H) exhibits better performance compared to both Vim-F and Vim-F(CF), and it is competitive with state-of-the-art ViT architectures.
5.3 Semantic Segmentation
Settings. Following Vim [22], we train UperNet [18] with our models on ADE20K dataset. In training, we employ AdamW with a weight decay of 0.01, and a total batch size of 16 to optimize models. The employed training schedule uses an initial learning rate of $6\times 10^{-5}$, linear learning rate decay, a linear warmup of 1500 iterations, and a total training of 160K iterations.
Table 5: Results of semantic segmentation on ADE20K.
Backbone Params mIoU
Vim-Ti [22] 34M 40.1
Vim-Ti-F (ours) 34M 40.4
ResNet-50 [4] 67M 40.7
Vim-S [22] 57M 43.9
Vim-S-F (ours) 58M 44.2
Vim-F(CF) (ours) 55M 44.4
Swin-T [10] 60M 44.5
Vim-S-F(H) (ours) 47M 44.8
Results. The results are presented in Table 5. Similar to the experimental results on MSCOCO, Vim-F and its variants demonstrate superior performance to Vim [22], ResNet [4], and Swin[10] in the semantic segmentation task, which further confirms the effectiveness of our work.
6 Conclusion
@@
One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data. Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships. We observe that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field. We adapt a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains. Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps. We also introduce additional local correlations during patch embedding by using overlapping convolutions. Vim-F(H) enhances the performance of Vim-F by introducing linear attention, which mitigates the sequence order sensitivity of the Mamba encoder. We constructed a proof-of-concept network, Vim-F(CF), and found that even after removing the convolutional layers from Vim-F, the model can still effectively model visual information. Extensive experiments show that our work effectively boosts Vim’s performance. Vim-F, Vim-F(H), and Vim-F(CF) achieve results competitive with advanced ViM models.